A sinusoidal wave traveling in the positive x direction has an amplitude of 15.0 cm, a wavelength of 40.0 cm, and a frequency of 8.00 Hz. The vertical position of an element of the medium at t = 0 and x = 0 is also 15.0 cm, Determine the phase constant , and write a general expression for the wave function 


1
Expert's answer
2022-03-23T07:15:16-0400

Given:

"A=15.0\:\rm cm"

"\lambda=40.0\:\rm cm"

"f=8.00\:\rm Hz"

"y(0,0)=15.0\:\rm cm"


The general expression for the wave function

"y(x,t)=A\cos\left(\frac{2\pi}{\lambda}x-2\pi ft+\phi_0\right)"


"y(x,t)=15.0\cos\left(\frac{2\pi}{40.0}x-2\pi*8.00t+\phi_0\right)"

"y(x,t)=15.0\cos\left(0.157x-20.2t+\phi_0\right)"

The initial condition gives

"y(0,0)=15.0\cos\left(\phi_0\right)=15.0"

Hence

"\cos\left(\phi_0\right)=1,\quad \phi_0=0"

Finally

"y(x,t)=15.0\cos\left(0.157x-20.2t\right)"